Find the solution of the initial-value problem and determine its interval of existence (i.e domain of the resulting function y). (i) dy dt = 1+t2 t , (ii) (t + 1)dy y(t = 1) = 0. dt = 1−y, y(t = 0) = 3. [Hint: Let A = −(±e−C).]
1
Expert's answer
2020-06-29T19:19:31-0400

i).Given,

"\frac{dy}{dt}=1+t^2"

Initial condition is not given,thus we are solving the general solution.


"dy=(1+t^2)dt\\
\implies \int dy=\int(1+t^2)dt\\\implies y=t+\frac{t^3}{3}+C"

Where, "C" is constant.

Clearly, "y" is a family of cubic polynomial in "t" ,thus domain is "\mathbb{R}"


ii).Question is badly stated.